OpenStudy (anonymous):

ok sorry for the stupid question but how do you find the area of a triangle?

OpenStudy (anonymous):

A = 1/2 b*h or A= b*h/2

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

your very welcome :)